OpenStudy (anonymous):

The radioactive substance uranium-240 has a half-life of 14 hours. The amount A(t) of a sample of uranium-240 remaining (in grams) after t hours is given by the following exponential function. Find the initial amount in the sample and the amount remaining after hours. Round your answers to the nearest gram as necessary. Initial Amount Amount after 40 hours

OpenStudy (anonymous):

\[A(t)=2800\left( \frac{ 1 }{ 2 } \right)^\frac{ t }{ 14 }\]

OpenStudy (anonymous):

@ikram002p @jigglypuff314 @phi can one of you help me with this?

OpenStudy (anonymous):

@jim_thompson5910 @whpalmer4 can one of you show me how to do this?

OpenStudy (whpalmer4):

Initial amount is simply \[A(0) = 2800(\frac{1}{2})^{\frac{0}{14}} = \] Amount at 40 hours is \[A(40) = 2800(\frac{1}{2})^{\frac{40}{14}} = \]